# Dynamic 3D Gaussian Fields for Urban Areas

### v1

@@ 4.1 Comparison to State-of-the-Art
4.2 Ablation Studies
We verify our design choices in both the multi- and single-sequence setting. For a fair comparison, we set the global maximum of 3D Gaussians to 8 and 4.1 million, respectively. We perform these ablation studies on the residential split of [17]. We use the full overlap in the multi-sequence setting, while using a single sequence of this split for the single-sequence setting. In Table 4a, we verify the components that are not specific to the multi-sequence setting. In particular, we show that our approach to modeling scene dynamics is highly effective, evident from the large disparity in performance between the static and the dynamic variants. Next, we show that modeling appearance with a neural field is on par with the proposed solution in [18], while being more memory efficient. In particular, when modeling view-dependent color as a per-Gaussian attribute as in [18] the model uses 8.6 GB of GPU memory during training, while it uses only 4.5 GB with fixed-size neural fields. Similarly, storing the parameters of the former takes 922 MB, while the latter takes only 203 MB. Note that this disparity increases with the number of 3D Gaussians per scene. Finally, we achieve the best performance when adding the generated 3D Gaussian background.
@@ 4.2 Ablation Studies
Figure 4: Qualitative results on Waymo Open [23]. We show a sequence of evaluation views synthesized by our model (top-left to bottom-right). As the woman gets out of the car, we successfully model her articulated motion and changing body poses.
@@ 5 Conclusion
Limitations. While 4DGF improves novel view synthesis in dynamic urban areas, the challenging nature of the problem leaves room for further exploration. Although we model scene dynamics, appearance, and geometry variations, other factors influence image renderings in real-world captures. First, in-the-wild captures often exhibit distortions caused by the physical image formation process. Thus, modeling phenomena like rolling shutter, white balance, motion and defocus blur, and chromatic aberrations is necessary to avoid reconstruction artifacts. Second, the assumption of a pinhole camera model in [18] persists in our work and thus our method falls short of modeling more complex camera models like equirectangular cameras which may be suboptimal for certain capturing settings.
@@ A.2 Method
In addition to the adjustments described above and inspired by recent findings [60], we adapt the criterion of densification during ADC. In particular, Kerbl et al. [18] use the average absolute value of positional gradient $\frac{\partial\mathcal{L}}{\boldsymbol{\mu}_{k}}$ across multiple iterations. The positional gradient of a projected 3D Gaussian is the sum of the positional gradients across the pixels it covers:
$\frac{\partial\mathcal{L}}{\boldsymbol{\mu}_{k}}=\sum_{i}\frac{\partial\mathcal{L}}{\partial\mathbf{p}_{i}}\frac{\partial\mathbf{p}_{i}}{\partial\boldsymbol{\mu}_{k}}\,.$ (15)
@@ A.3 Experiments
In our multi-sequence experiments in Table 1 and Table 4, we train our model on 8 NVIDIA A100 40GB GPUs for 125,000 steps, taking approximately 2.5 days. In our single-sequence experiments, we train our model on a single RTX 4090 GPU for several hours. On Waymo Open, we train our model for 60,000 steps while for KITTI and VKITTI2 we train the model for 30,000 steps. For our singl-sequence experiments in Table 4 we use a schedule of 100,000 steps. We chose a longer schedule for Waymo Open and Argoverse 2 since the scenes are more complex and contain about 510×5-10$\times$5 - 10 × as many images as the sequences in KITTI and VKITTI2.
@@
Additional ablation studies. In Table 6a, we show that while our approach benefits from high-quality 3D bounding boxes, it is robust to noise and achieves a high view synthesis quality even with noisy predictions acquired from a 3D tracking algorithm [87]. In Table 6b, we demonstrate that the deformation head yields a small, albeit noticeable improvement in quantitative rendering results. This corroborates the utility of deformation head $\chi$ beyond the qualitative examples shown in Figures 4 and 8. Note that the threshold to distinguish between dynamic and static areas is 1m/s following [76] so that some instances like slow-moving pedestrians will be classified as static. Also, since non-rigid entities usually cover only a small portion of the scene, expected improvements are inherently small.
@@
Figure 8: Additional qualitative results on Waymo Open [23]. We show a sequence of evaluation views synthesized by our model (top to bottom). We see two pedestrians on the left and right being faithfully modeled across varying body poses while the right one also carries a stroller.

### v2

@@ 4.3 Runtime analysis
Dataset Argoverse 2 [81] Waymo Open [23] Mean (ms) Percentage (%)
Resolution 1550 $\times$ 2048 640 $\times$ 960
Avg. number of 3D Gaussians 8.02M 2.75M
1. Scene graph evaluation: retrieve $\omega$, $[\mathbf{R}|\mathbf{t}]$, 3D Gaussians at $(s,t)$ 38.5 13.0 25.75 52.4
2. Scene composition: apply $[\mathbf{R}|\mathbf{t}]$ to 3D Gaussians 2.3 1.5 1.90 3.9
3. 3D Gaussian projection 2.0 3.5 2.75 5.6
4. Query neural fields $\phi$ and $\psi$ 9.5 3.6 6.55 13.3
5. Rasterization 21.3 3.0 12.15 24.8
Total 73.6 24.6 49.1 100
FPS 13.6 40.7 20.4 -
Table 6: Inference runtime analysis. We report the individual and average timings of our method’s components on two datasets. Overall, scene graph evaluation (1.) and rasterization (5.), dominate the runtime. While total runtime correlates with scene scale and image resolution, we achieve interactive frame rates on both datasets.
@@
Figure 4: Qualitative results on Waymo Open [23]. We show a sequence of evaluation views synthesized by our model (top-left to bottom-right). As the woman (marked with a red box) gets out of the car and walks away, we successfully model her articulated motion and changing body poses.
@@ 5 Conclusion
Limitations and future work. While 4DGF improves novel view synthesis in dynamic urban areas, the challenging nature of the problem leaves room for further exploration. Although we model scene dynamics, appearance, and geometry variations, other factors influence image renderings in real-world captures. First, in-the-wild captures often exhibit distortions caused by the physical image formation process. Therefore, modeling phenomena like rolling shutter, white balance, motion and defocus blur, or chromatic aberrations is necessary to avoid reconstruction artifacts. Second, the assumption of a pinhole camera model in [18] persists in our work. Thus, our method falls short of modeling more complex camera models like equirectangular cameras and other sensors such as LiDAR, which may be limiting for certain capturing or simulation settings.
@@ A.2 Method
In addition to the adjustments described above and inspired by recent findings [60], we adapt the criterion of densification during ADC. In particular, Kerbl et al. [18] use the average absolute value of positional gradient $\frac{\partial\mathcal{L}}{\partial\boldsymbol{\mu}_{k}}$ across multiple iterations. The positional gradient of a projected 3D Gaussian is the sum of the positional gradients across the pixels it covers:
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\mu}_{k}}=\sum_{i}\frac{\partial\mathcal{L}}{\partial\mathbf{p}_{i}}\frac{\partial\mathbf{p}_{i}}{\partial\boldsymbol{\mu}_{k}}\,.$ (15)
@@ A.3 Experiments
In our multi-sequence experiments in Table 1 and Table 5, we train our model on 8 NVIDIA A100 40GB GPUs for 125,000 steps, taking approximately 2.5 days. In our single-sequence experiments, we train our model on a single RTX 4090 GPU for several hours. On Waymo Open, we train our model for 60,000 steps while for KITTI and VKITTI2 we train the model for 30,000 steps. For our single-sequence experiments in Table 5 we use a schedule of 100,000 steps. We chose a longer schedule for Waymo Open and Argoverse 2 since the scenes are more complex and contain about 510×5-10$\times$5 - 10 × as many images as the sequences in KITTI and VKITTI2.
@@
Additional ablation studies. In Table 8a, we show that while our approach benefits from high-quality 3D bounding boxes, it is robust to noise and achieves a high view synthesis quality even with noisy predictions acquired from a 3D tracking algorithm [88]. In Table 8b, we demonstrate that the deformation head yields a small, albeit noticeable improvement in quantitative rendering results. This corroborates the utility of deformation head $\chi$ beyond the qualitative examples shown in Figures 4 and 10. Note that the threshold to distinguish between dynamic and static areas is 1m/s following [76] so that some instances like slow-moving pedestrians will be classified as static. Also, since non-rigid entities usually cover only a small portion of the scene, expected improvements are inherently small.
@@
Figure 7: Runtime comparison of neural fields vs. spherical harmonics. We compare the runtime of querying neural fields $\phi$ and $\psi$ to a spherical harmonics function of degree 3. We report time-per-query in nanoseconds.
Figure 8: Histogram of mean 3D Gaussian scales. We use our model trained on Argoverse 2 (residential split). Both axes are in logarithmic scale. The vast majority of 3D Gaussians have a small scale, while there are a few outliers with huge scales. The scene is approximately within [-1, 1].
Additional runtime and model analysis. In Figure 8, we provide a comparison in terms of time per query of neural fields of different sizes (equivalent to the sizes of our $\phi$ and $\psi$) versus querying a spherical harmonics function of degree 3 as used in 3DGS [18]. We observe that the SH function is approx. 6 to 8$\times$ faster to evaluate than the neural fields. However, we show that this does not lead to a critical increase in overall runtime in Table 6. Additionally, we note that the SH function is limited in representation capacity: It is not capable of handling varying appearance across input sequences (weather, time of day, season), transient geometry (construction sites, tree leaves), articulated motion of dynamic objects (pedestrians, cyclists), and large-scale scenes with several millions of 3D Gaussians due to memory constraints. Thus, we emphasize that the use of neural fields does not merely improve memory footprint, but enables applications that 3DGS [18] is not capable of modeling.
In Figure 8, provide an analysis on the distribution of 3D Gaussian scales across a large-scale urban scene reconstruction. We notice that while the vast majority of 3D Gaussians is small ($<0.001$ mean scale), there is a small number of very large 3D Gaussians ($>1.0$ mean scale) compared to the scene bounds that are approximately within [-1.0, 1.0]. This can lead to fog-like artifacts in free-viewpoint renderings. We note that this can be mitigated using a regularization term on the 3D Gaussian scales, however, we did not observe an quantitative improvement in view synthesis metrics and thus we did not include it in the experiments.
@@
Figure 10: Additional qualitative results on Waymo Open [23]. We show a sequence of evaluation views synthesized by our model (top to bottom). We see two pedestrians on the left and right being faithfully modeled across varying body poses while also carrying objects such as a stroller (right) or a shopping bag (left).